OpenStudy (sasogeek):

Evaluate \[\large \lim_{x \rightarrow 2}[ln(4x+1)] \]

OpenStudy (rogue):

Just plug in 2 for x and simplify.\[\large \lim_{x \rightarrow 2}[\ln(4x+1)] = \ln(4*2+1) = \ln 9 = 2\ln3\]